OpenStudy (anonymous):

implify. fourthroot(a^6) - (2a)(sqrt a). click to see better format

OpenStudy (anonymous):

\[\sqrt[4]{?x^6}-2x \sqrt{x}\]

OpenStudy (anonymous):

oops. \[\sqrt[4]{x^6}-2x \sqrt{x}\]

OpenStudy (anonymous):

\[\sqrt[4]{a}^{6}=a ^{6\times1/4}\] \[a ^{3\div2}-2a \sqrt{a}\] \[\sqrt[2]{a}^{3}-2a \sqrt{a}\]

OpenStudy (anonymous):

isnt it 4 x 1/6

OpenStudy (anonymous):

instead of 6 x 1/4

OpenStudy (anonymous):

hey saifoo i need you to be my personal tutor for tonite. i got a big test tomorrow! could ya? please.?

OpenStudy (saifoo.khan):

well, i can help.. But im not very good with these. but still, i can help!

OpenStudy (anonymous):

thanks so much. ok. so you dont know this one?

OpenStudy (saifoo.khan):

i can do this.

OpenStudy (anonymous):

yeah its pretty hard. umm how about..

OpenStudy (anonymous):

\[(2/x)-(2/(2x+1))=1/x+1\]

OpenStudy (anonymous):

find x ^^^

OpenStudy (saifoo.khan):

brb.

OpenStudy (saifoo.khan):

5mins!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

it would be WAY easier if i opened up a classroom. cuz i got like 10 problems. but either way is fine. im still getting help.

OpenStudy (saifoo.khan):

http://www.twiddla.com/579480

OpenStudy (saifoo.khan):

im already teaching someone! u may join!

OpenStudy (anonymous):

what happened? you got kicked?

OpenStudy (anonymous):

"isnt it 4 x 1/6" No, here are the rules: a^b * a^c = a^{b+c} a^b/a^c =a^{b-c} (a^b)^c = a^{bc} You have x^6 and you want the fourth root of it ie the quarter power, so (x^6)^(1/4) = x^(6*1/4) = x^(3/2) The other term is 2x (think 2x^1) multiplied by the half power 2xx^(1/2) = 2x^(3/2) so finally u have x^(3/2) - 2x^(3/2) = -x^(3/2) which you can check at Wolfram http://www.wolframalpha.com/input/?i=%28x^6%29^%281%2F4%29+-+2x*sqrt%28x%29